OpenStudy (anonymous):

a-4a=2a+1-5a (A number,No Solution,All real numbers,or Many solutions?)

OpenStudy (anonymous):

a - 4 a = 3a 3a = 2a+ 1 - 5a bring the -5a over changes to +5a 5a + 3a = 2a + 1 8a = 2a + 1 bring the 2a over changes to -2a 2a - 8a = 1 6a = 1 divide both sides by 6 6a/6 = 1/6 a = 1/6